### v4

# Repository-level Context Code Translation Benchmark Targeting Rust

Abstract.
Recent advancements in large language models (LLMs) have demonstrated impressive capabilities in code translation, typically evaluated using benchmarks like CodeTransOcean. However, these benchmarks fail to capture real-world complexities by focusing primarily on simple function-level translations and overlooking repository-level context (e.g., dependencies). Moreover, LLMs’ effectiveness in translating to newer, low-resource languages like Rust remains largely underexplored.
To address this gap, we introduce RustRepoTrans, the first repository-level context code translation benchmark, comprising 375 tasks translating into Rust from C, Java, and Python. Using this benchmark, we evaluate four state-of-the-art LLMs, analyzing their errors to assess limitations in complex translation scenarios. Among them, Claude-3.5 performs best with 43.5% Pass@1, excelling in both basic functionality and additional translation abilities, such as noise robustness and syntactical difference identification. However, even Claude-3.5 experiences a 30.8% performance drop (Pass@1 from 74.3% to 43.5%) when handling repository-level context compared to previous benchmarks without such context. We also find that LLMs struggle with language differences in complex tasks, and dependencies further increase translation difficulty.
copyright: acmlicensedjournalyear: 2025
1. Introduction
Code translation, or code migration, is the process of converting a software project from one programming language to another (Pan et al., 2023), often driven by the need to adapt to different runtime environments, improve performance, or enhance security (nee, 2018, 2020b, 2020c, 2020d, 2020a; Thönes, 2015). The rise of languages like Rust (rus, 2024) and Cangjie (can, 2025) has increased demand for translation. However, migrating legacy code remains challenging due to semantic mismatches, differences in standard libraries, and the need to maintain functional equivalence (Mateus et al., 2023; Kontogiannis et al., 2010).
Large language models (LLMs) have shown promise in translating widely used languages like Java and Python (Pan et al., 2024, 2023), but their effectiveness in translating to newer, lower-resource languages like Rust has not been thoroughly explored. Rust, an emerging language known for its safety guarantees, is increasingly in demand for translating projects from other languages to enhance memory safety and reliability (Security, 2024). However, its distinct features and limited training data pose challenges for automated translation, potentially limiting existing models’ performance (rus, 2024).
Existing code translation benchmarks, while covering Rust to some extent, have several limitations in capturing real-world scenarios. Most of them focus on function-level translation pairs (Roziere et al., 2020; Lu et al., 2021; Zheng et al., 2023; Jiao et al., 2023), often translating self-contained functions with minimal dependencies (Puri et al., 2021; Yan et al., 2023; Zhu et al., 2022a). Even file-level benchmarks only provide limited context for making isolated functions executable. Furthermore, many datasets are sourced from online coding platforms or artificially created examples, which differ greatly from real-world development practices (Puri et al., 2021; Yan et al., 2023; Zhu et al., 2022a; Roziere et al., 2020; Lu et al., 2021; Zheng et al., 2023; Jiao et al., 2023). In contrast, real-world translation occurs at the repository level, where developers incrementally translate code while preserving the project’s architecture. This involves managing complex dependencies, cross-file references, and inconsistencies between large-scale projects, aspects that current benchmarks fail to capture.
To bridge this gap, we introduce RustRepoTrans, the first benchmark to evaluate code translation with repository-level context, specifically targeting Rust. Unlike function-level benchmarks that assess translation in isolation, RustRepoTrans captures the complexities of real-world migration by incorporating repository-level context, including dependencies, cross-file interactions, and architectural constraints. These elements are crucial for ensuring that translated code integrates seamlessly into existing projects rather than functioning as isolated snippets. Our benchmark comprises 375 curated translation tasks derived from real-world projects spanning C, Java, and Python to Rust, providing a more faithful assessment of LLMs’ translation capabilities. While we focus on Rust due to its growing adoption in systems programming and the inherent challenges of translating into it, our methodology is generalizable to other programming languages, paving the way for broader repository-level evaluation.
Building on RustRepoTrans, we assess state-of-the-art LLMs in translating code with complex repository-level context, analyzing challenges beyond existing benchmarks. We identify key factors affecting task difficulty and common errors, while also evaluating noise robustness, syntactical difference handling, and code simplicity during translation.
Our experiments show that LLMs struggle with repository-level code translation, with compilation errors reaching 94.8%, exposing the gap between current evaluations and real-world performance. Claude-3.5 performs best with 43.5% Pass@1, excelling in basic functionality and additional translation abilities like noise robustness and syntactical difference identification. However, even Claude-3.5 sees a 30.8% drop (Pass@1 from 74.3% to 43.5%) when handling repository-level context. Dependency-related errors, including function and variable resolution issues, account for 61.9% of failures, highlighting the challenge of interconnected code. We also find that code complexity correlates with translation effectiveness. These insights reveal LLMs’ limitations in complex code migration, guiding future improvements in model design and evaluation.
The contributions of this paper are as follows. Data and code are publicly available at  (rep, 2024).
  • The first repository-level code translation benchmark, RustRepoTrans, with 375 tasks from C, Java, and Python to Rust, is introduced for realistic evaluation.
  • Four state-of-the-art LLMs are evaluated on this benchmark, assessing their performance in real-world translation tasks.
  • LLM errors are categorized into 10 types, revealing limitations, especially in handling dependencies.
2. Related Work
2.1. Code Translation
Code translation enhances software interoperability and facilitates legacy system updates, boosting productivity and reducing manual costs (Lu et al., 2021; Puri et al., 2021). Early methods relied on rule-based systems with handcrafted syntactic and statistical machine translation (SMT) models, improving accuracy by mapping frequent code patterns (Nguyen et al., 2013; Chen et al., 2018). Neural models, including RNNs, improved learning from data but struggled with complex syntax and idioms (Dong and Lapata, 2016; Yin and Neubig, 2017, 2018).
Despite advancements with LLMs, translation errors and adaptability issues persist (Zhu et al., 2022a; Rithy et al., 2022). Jiao et al. (Jiao et al., 2023) highlighted LLMs’ difficulty with complex translations, while Pan et al. (Pan et al., 2024) showed that context-rich prompts improve reliability by 5.5%. Yang et al. (Yang et al., 2024) and Macedo et al. (Macedo et al., 2024a) further refined prompt engineering for accuracy and consistency.
Our work introduces RustRepoTrans, which enables a more realistic evaluation of LLMs by incorporating repository-level dependencies and context. This approach allows us to better assess existing code translation techniques in handling complex code translation tasks.
2.2. Code Translation Benchmarks
Recent advances in neural code translation models have been fueled by high-quality, large-scale parallel datasets. However, existing datasets exhibit significant limitations in programming language coverage, scale, and balance, particularly for Rust. Many datasets focus narrowly on a few languages, such as the early studies between Java and C (Nguyen et al., 2013) and more recent work like CodeTrans (Lu et al., 2021) and AVATAR (Ahmad et al., 2021), which primarily cover Java and Python.
To address these gaps, some datasets, such as CoST (Zhu et al., 2022b) and CodeNet (Puri et al., 2021), have aimed to enhance language diversity and quality. However, issues like erroneous solutions and limited support for Rust persist. The latest benchmark, CodeTransOcean (Yan et al., 2023) and XCodeEval (Khan et al., 2024a), expands coverage to low-resource languages but still lacks adequate support and error analysis for Rust, hindering comprehensive research in this area.
Furthermore, many existing benchmarks rely on data from question-and-answer websites, which do not reflect real-world software development contexts. These datasets fail to capture the complex dependencies and architectural considerations inherent in actual projects, where code translation often requires maintaining the original architecture and addressing intricate cross-file dependencies. This disconnect underscores the need for more representative benchmarks like RustRepoTrans, which aim to provide a more realistic assessment of code translation capabilities.
RustRepoTrans is the first repository-level context code translation benchmark focusing on Rust, designed to account for complex dependencies and derived from real-world projects, distinguishing it from previous benchmarks.
3. RustRepoTrans Benchmark
We introduce the benchmark format, construction method, and the resulting RustRepoTrans.
3.1. Benchmark Format
Figure 1. RustRepoTrans Format Example
Each task in RustRepoTrans consists of a pair of functions with their dependencies and test cases, formatted as <source function, target function, target function dependencies, target function test cases> (Fig. 1). The function pairs represent functionally equivalent code snippets from source and target languages, with dependencies relevant to target function such as functions, data types, variables, and libraries. LLMs use the source function, target function signature, and associated dependencies to generate the target function, which is verified for correctness using the test cases.
3.2. Benchmark Construction Method
The construction process is divided into two parts: Functionally Equivalent Code Pairs Extraction and Dependency Extraction. This approach allows us to obtain functionally equivalent code pairs along with their corresponding dependencies and test cases from real open-source projects.
Functionally Equivalent Code Pairs Extraction In this part, we focus on extracting functionally equivalent code pairs (source function and target function), with the target function written in Rust, from GitHub projects. The process involves five steps, detailed as follows.
Migration Project Selection. In this step, we select projects that have been rewritten in Rust from other languages, specifically C, Java, and Python, due to their popularity and likelihood of having such rewritten versions. These projects are expected to contain functionally equivalent function pairs between their Rust and original language versions. We identify suitable projects by searching GitHub for terms like “Rust version” and “implemented in Rust”. focusing on larger projects (applying qualifier “size:>1000”) to ensure sufficient function pairs. After locating these projects, we verify their versions through documentation and code review. In this way, we identify pairs of source project and target project.
Functions Pools Extraction. In this step, we extract all functions from a pair of projects (source and target). For the Rust target project, we focus on candidate functions with associated test cases to ensure verifiability. We first identify all test cases and then extract the functions they cover by analyzing the external functions invoked. For the source project, we extract all implemented functions for comprehensive coverage. This extraction uses the static code analysis tool tree-sitter (tree sitter, 2023). As a result, we obtain two sets of functions: the source functions pool and the target functions pool with test cases.
Similarity-based Candidate Function Pair Extraction. In this step, we extract function pairs from the source and target function pools using a similarity-based approach. Developers often translate code at the function level while maintaining a similar structure across languages. Thus, equivalent pairs usually come from files with similar paths and function signatures (Zhong et al., 2010). For example, the Rust function $pbkdf2$ at $src/ecdh.rs$ corresponds to $PBKDF2$ in the C project at $src/ecdh\_support.c$. We use the BM25 algorithm (Trotman et al., 2014) to calculate the similarity for each target function, identifying the top 10 candidate source functions. This produces a list of Rust target functions with their top-10 source function candidates.
LLM-based Equivalent Function Pair Identification. In this step, we identify the most equivalent function for each Rust target function from its top-10 source function candidates using an LLM, leveraging their code implementation and contextual information (such as file paths). LLMs like GPT-4 excel in code understanding (Li et al., 2023). The prompt used for this identification is shown in Fig. 2. We employ GPT-4o due to its effective balance of efficiency and performance. If none of the candidates are functionally equivalent, the LLM is instructed to select "None."
Figure 2. Prompt for Identifying Equivalent Function Pairs
Manual Verification. Due to the limitations of LLMs, such as the potential for hallucinations, we manually verify the functionally equivalent pairs identified by the models to ensure their actual functional equivalence.
Dependencies Extraction. In this part, we complement the extracted pairs of functionally equivalent functions (source function, target function, test cases) with dependencies from the target projects. This approach adds a unique repository-level context for code translation, distinguishing real-world functions from those manually constructed or sourced from programming Q&A websites.
We classify dependencies into three categories: invocation function dependencies, invocation variable dependencies, and data type dependencies. To identify these, we perform static analysis on the entire project to extract custom functions, data types, and global variables. For the target function, we gather import statements, call function identifiers, variable dependencies, and data type dependencies. We match in-file dependencies and, using the import statements, match cross-file dependencies to compile a complete list for the function. Due to tree-sitter’s limitations, we manually reviewed the automatically extracted dependencies to correct any errors or omissions, ensuring each function has a complete and accurate set of dependencies.
3.3. Resulting Benchmark
This process constructs RustRepoTrans with 375 repository-level translation tasks. Table 1 compares it with existing benchmarks. Similar in size to prior datasets, each task includes a manually verified ground truth translation and unit tests, achieving over 90% test coverage, ensuring high quality.
RustRepoTrans has two key features that distinguish it from previous benchmarks: 1) it is the first code translation benchmark to consider repository-level dependencies; and 2) it specifically targets code translation to Rust in realistic programming scenarios. Next, we will discuss these features in more detail.
Table 1. Comparison of Different Benchmarks for Code Translation
Dataset Source Task Level #Tokens # Tasks Source Languages With Dependencies? Target Languages Include Rust? Unit Tests Included ? Golden Answer Verified?
CodeXGLUE(Lu et al., 2021) Lucune, POI, JGit, Antlr Function Level 42.3 1,000 Java, C#
XLCOST(Zhu et al., 2022a) G4G Program Level 202 901 C++, Java, C#, PHP, JavaScript, Python, C
TransCoder-test(Roziere et al., 2020) G4G Function Level 107.0 948 C++, Java, Python Partial
HumanEval-X(Zheng et al., 2023) HumanEval Program Level 97.7 164 C++, Java, Go, JavaScript, Python
G-TransEval(Jiao et al., 2023) HumanEval, G4G, .NET samples Function Level 95.1 400 C++, Java, C#, JavaScript, Python
CodeTransOcean(Yan et al., 2023) Rosetta Code Program Level 448.4 2,878* Java, C++, C#, PHP, Python, Go
RustRepoTrans Github Function Level 150.1 375 C, Java, Python
*Translation pairs in languages that could not be parsed by tree-sitter to count the number of tokens were filtered out.
Repository-level Dependency. RustRepoTrans focuses on code translation tasks with rich repository context, setting it apart from previous benchmarks. Unlike artificially constructed data or data from Q&A websites, real project data exhibits more complex dependency relationships, including function, data type, and variable dependencies, which are absent in existing datasets as shwon in Table 1, revealing that only RustRepoTrans includes these crucial elements. This inclusion makes RustRepoTrans a more realistic benchmark, suitable for evaluating LLMs that must account for intricate file-level interactions and contextual dependencies.
Rust Programming in Realistic Scenarios. RustRepoTrans and CodeTransOcean are the only benchmarks specifically targeting code translation to Rust, as listed in Table 1. However, RustRepoTrans is derived from GitHub projects, making it more reflective of real-world development than CodeTransOcean, which relies on data from programming competition websites.
4. Evaluation
Based on RustRepoTrans, we further investigate the performance of the state-of-the-art LLMs on repo-level code translation task for Rust. Specifically, we focus on the following RQs.
  • RQ1 (LLMs Performance): How do the state-of-the-art LLMs perform on RustRepoTrans in terms of translation effectiveness?
  • RQ2 (RustRepoTrans Effectiveness): How effectively does our new benchmark pose greater challenges in code translation compared to existing benchmarks?
  • RQ3 (Difficulty Analysis): What factors affect the difficulty of the translation tasks?
  • RQ4 (Failure Analysis): What types of errors do LLMs encounter on RustRepoTrans, and what factors contribute to these translation failures?
  • RQ5 (Key Capabilities): Beyond translation accuracy, what essential capabilities do LLMs demonstrate on RustRepoTrans, including noise robustness, syntactical difference identification, and code simplicity?
4.1. Experimental Setup
Model Selection. We selected four state-of-the-art LLMs that have been widely studied in recent code translation research or were recently released and have demonstrated strong performance on benchmark leaderboards (e.g.,SuperCLUE (Xu et al., 2023)). Table 2 provides an overview of the four LLMs used in our experiments, along with their release dates (Column “Time”) and model sizes (Column “Size”). As shown in Table 2, our study includes both open-source and closed-source LLMs, as well as models designed for general-purpose and code-specific applications. Due to resource constraints, we selected open-source models with sizes below 20B parameters. Note that our primary contribution is a new benchmark designed to provide a more challenging and realistic evaluation of LLMs’ code translation capabilities, closely simulating real-world development scenarios. Expanding the scope to include a larger and more diverse set of models is reserved for future work.
Implementation Details. For open-source models, we obtained and executed the released versions from their official repositories based on provided documentation with greedy sampling (Chen et al., 2015) as our generation strategy. All evaluations were conducted on an NVIDIA A800 80GB GPU. For closed-source LLMs, we accessed each model through its respective API interface (OpenAI, 2024; Claude, 2024) as of September 2024. To achieve results similar to greedy decoding, we set the “temperature” hyperparameter to 00.
4.2. RQ1: LLMs Performance
We evaluated four selected LLMs on RustRepoTrans to assess their translation accuracy and error-correction abilities for complex, repository-level code translation tasks targeting Rust.
4.2.1. Design
The evaluation involved testing each model on RustRepoTrans using controlled prompts, with output correctness assessed through specific test cases.
Evaluation Process. Each selected LLM was tasked with translating code into Rust, specifically focusing on 375 tasks. For each translation task, a corresponding set of test cases was employed to evaluate the correctness of the generated outputs. To ensure a fair comparison, the same prompt was used for each model, carefully designed based on established best practices in code translation tasks, as illustrated in Fig. 3. These best practices have been shown to enhance translation accuracy in similar studies (Macedo et al., 2024b; Khan et al., 2024b; Pan et al., 2024; Yang et al., 2024). The prompt includes elements such as Instruction for translation and Translation’s required information(including source code, target function signature and target function dependencies)
Table 2. Studied LLMs
Model Type Model Company
Open-source
Time Size
General LLM Claude-3.5 (Anthropic, 2024) Anthropic 2024.6 -
GPT-4 (OpenAI, 2023) OpenAI 2023.6 -
Llama-3.1-8B (Meta, 2024) Meta 2024.7 8B
Code LLM DeepSeekCoderV2-16B (DeepSeek, 2024) DeepSeek 2024.6 16B
Given that existing LLM-based code translation research often incorporates feedback on translation error messages to enhance performance (Yang et al., 2024; Pan et al., 2024), further experiments were conducted to understand the extent to which errors in the generated code samples can be corrected by LLMs with feedback.
It is important to note that the goal is to evaluate the LLMs’ intrinsic ability to self-correct translation errors, as well as to explore how many translation errors in RustRepoTrans can be automatically corrected by LLMs. Therefore, in addition to self-debugging, the best-performing model on RustRepoTrans, Claude-3.5, was also used to debug the translation outputs of all models. The debugging prompt (Prompt-Fix) shown in Fig. 3 includes Instruction for previous translation, Instruction for debugging, Incorrect translation and error details and Translation’s required information, which are adapted from previous work (Pan et al., 2024; Yang et al., 2024).
Metrics. Two key metrics were utilized to measure performance:
  • Pass@k: This widely-used metric (Chen et al., 2021) calculates the percentage of tasks correctly solved based on $k$ generated code samples per task, as expressed in Equation 1. A task is considered solved if at least one of the generated code samples passes all the corresponding test cases. Following recent research (Pan et al., 2024), the focus was on calculating the Pass@1 metric, where $k=1$. This approach reflects the model’s ability to produce a correct translation on the first attempt.
[…]
  • DSR@k: The DSR@k (Debugging Success Rate@k) metric, proposed in previous work (Yan et al., 2023), evaluates whether the generated code successfully executes and produces the expected results (i.e.,passing all test cases) within $k$ rounds of debugging, as expressed in Equation 2. $S(i,k)=1$ if the $i^{th}$ code sample succeeds within $k$ attempts; otherwise $S(i,k)=0$. In this study, DSR@k metrics were calculated with $k=1$, allowing for one debugging attempt.
[…]
Figure 3. Translating Prompt and Debugging Prompt
4.2.2. Results
Fig. 4 presents the Pass@1 and DSR@1 performance of each LLM on RustRepoTrans. This benchmark challenges models with tasks that include complex dependencies, offering insights into each model’s initial translation accuracy and post-debugging improvement potential.
Initial Translation Performance (Pass@1). In the initial code translation task, Claude-3.5 substantially outperformed the other LLMs, achieving 14.2% higher accuracy than GPT-4, 26.7% higher than DeepSeekCoderV2-16B, and 30.2% higher than Llama-3.1-8B on the Pass@1 metric. This indicates Claude-3.5’s comparative strength in handling repository-level code translation to Rust. However, even Claude-3.5 achieved only a 45.3% Pass@1 score, emphasizing the difficulty of RustRepoTrans and the challenges inherent in translating code with complex dependencies.
Post-debugging Performance. After a single round of self-debugging, each model’s DSR@1 score showed significant improvement, confirming previous research that LLMs can effectively leverage compiler feedback for code translation accuracy (Yan et al., 2023; Yang et al., 2024). However, improvements from debugging by Claude-3.5 surpassed those from self-debugging, with Llama-3.1-8B showing only a 5.6% increase from self-debugging compared to a 24.8% improvement from Claude-3.5. This underscores Claude-3.5’s superior debugging capabilities in complex code translation.
Post-debugging performance rankings revealed Claude-3.5 leading at 56.5%, followed by GPT-4 at 44.5%. Notably, Llama-3.1-8B (38.1%) surpassed DeepSeekCoderV2-16B (36.0%), indicating its initial errors were more correctable. Despite Claude-3.5 ’s strong performance, 44.5% of its cases remained unresolved, highlighting the inherent challenges of self-correction in intricate translation tasks. In terms of relative improvement, Llama-3.1-8B exhibited the highest percentage gain at 186%, indicating its initial errors were highly correctable. Conversely, Claude-3.5 showed the smallest improvement at 13.0%, suggesting its initial outputs were closer to correct execution. Overall, even after self-debugging, Claude-3.5 maintained its leading position, emphasizing its robust capabilities in repository-level code translation to Rust.
Figure 4. Pass@1, DSR@1(self-debugging) and DSR@1(debug by Claude-3.5) on RustRepoTrans
4.2.3. Summary.
The results reveal that RustRepoTrans effectively challenges LLMs in complex code translation, as even the best model, Claude-3.5, only achieves a Pass@1 of 45.3%, and DSR@1 of 56.5%, highlighting the limitations of LLMs.
4.3. RQ2: RustRepoTrans Effectiveness
To assess our benchmark’s effectiveness, a comparative analysis was conducted based on prior literature.
4.3.1. Design.
We evaluated performance on RustRepoTrans and a prior Rust translation benchmark without repository-level context, demonstrating RustRepoTrans’s effectiveness in assessing models for complex, real-world translation scenarios. Our analysis is based on CodeTransOcean (Yan et al., 2023), which includes translation tasks across 45 programming languages, including Rust, without repository-level dependencies. To ensure a fair and efficient comparison, we constructed a subset of CodeTransOcean by selecting translation pairs where Rust is the target language and C, Java, or Python is the source language, removing those with non-compilable Rust code. From the remaining pairs, we randomly sampled 300 (100 per language) to match our dataset’s scale for comparative experiments.
Claude-3.5 was selected for its top performance in RustRepoTrans. It generated one translation per task, evaluated using corresponding tests. We computed the average Pass@1 across source languages and compared it with RQ1. Prompts followed those in Fig. 3 but excluded dependency-related instructions. Hyperparameters were the same as in RQ1, with the temperature set to zero.
4.3.2. Results
The results presented in Table 3 reveal significant differences in the Pass@1 scores of Claude-3.5 across various datasets. Notably, while Claude-3.5 achieves high accuracy rates—ranging from 67.0% to 80.0%—on established benchmarks, its performance drastically drops when evaluated on RustRepoTrans.
Specifically, the Pass@1 scores for RustRepoTrans are significantly lower, with rates of 48.3% for C to Rust, 39.3% for Java to Rust, and 41.7% for Python to Rust translations, averaging 42.5%. This indicates a decline of approximately 25.3% to 36.7% compared to the higher scores observed on CodeTransOcean. This notable decrease highlights the increased complexity and unique challenges posed by RustRepoTrans, reinforcing the need for models to effectively manage repository-level dependencies and navigate intricate code structures in real-world.
Table 3. Pass@1 Comparison of Claude-3.5 in C, Java, and Python to Rust on CodeTransOcean and RustRepoTrans
Source Language Dataset Pass@1(Claude-3.5)
C CodeTransOcean 80.0%
RustRepoTrans 48.3%(↓31.7%)
Java CodeTransOcean 76.0%
RustRepoTrans 39.3%(↓36.7%)
Python CodeTransOcean 67.0%
RustRepoTrans 41.7%(↓25.3%)
4.3.3. Summary
Our benchmark’s repository-level context poses greater challenges for LLMs, reducing Claude-3.5’s Pass@1 from 74.3% to 43.5% (a 30.8% drop) in average. This underscores the benchmark’s effectiveness in evaluating real-world code translation and the need for models to better handle project dependencies.
4.4. RQ3: Difficulty Analysis
To understand the challenges that RustRepoTrans presents to LLMs and how they contribute to performance drops, we analyze the code translation results of LLMs across different tasks.
4.4.1. Design
We analyzed both successfully and unsuccessfully translated tasks using Claude’s results, chosen for its superior performance. Our goal was to identify the types of tasks where it succeeds or fails. Based on our dataset characteristics, we focused on three key factors that influence translation difficulty: the number of dependencies, and the lines of code in both the source and target languages. We conducted statistical analyses of these factors to assess their impact on translation success.
4.4.2. Results
Fig. 5 illustrates the distribution of successful and unsuccessful tasks with Claude based on dependencies and source/target code lines. Successful translations are concentrated on the left side, indicating fewer dependencies and shorter code lines. In contrast, unsuccessfully translated pairs show a wider distribution. Specifically, failed translations have, on average, 138.7% more dependencies, 134.6% more lines of code in the target language, and 108.4% more lines of code in the source language compared to successful ones. This suggests that failed function pairs present more challenging translation tasks, resulting in lower success rates.
Figure 5. Success and Failure Distribution of Claude by Dependencies and Source/Target Code Lines
4.4.3. Conclusion
Higher dependencies and longer code lines challenge LLMs, resulting in lower translation success rates.
4.5. RQ4: Failure Analysis
To understand the errors causing performance drops in LLMs during complex Rust code translation, we analyzed the erroneous results generated by the models.
4.5.1. Design
The analysis comprises two parts: an automatic error outcome analysis and a manual error causes analysis.
Automatic Error Outcome Analysis. We collected all unsuccessful translations from LLMs in RQ1 (Section 4.2), resulting in 1,114 code samples that failed to pass all test cases. To identify where LLMs struggle in translating code to Rust, we conducted an automated analysis following the methodology of Pan et al. (Pan et al., 2024), categorizing the unsuccessful translations into four error outcomes: compilation errors (code fails to compile), runtime errors (code compiles but encounters exceptions), functional errors (code executes successfully but fails test cases), and non-terminating execution (code runs indefinitely or waits for input).
Manual Error Causes Analysis. We further conducted a manual analysis of the error causes for the unsuccessful translations with compilation errors (1,056 code samples), which accounted for the vast majority, 94.8%, of the failures as shown in Figure 4 (see Section 4.5.2). Using an open coding approach (error codes, 2024), we examined the error messages for each failing result and classified each compilation error into specific categories. If an error could not fit into an existing category, we created a new one or adjusted the definition of an existing category to include it, followed by revising and reannotating all relevant samples.
Notably, the same type of error could belong to different categories depending on the context. This iterative annotation process involved collaborative discussions to refine the categories and ensure consistent classification. We regularly reviewed and reorganized the categories to maintain clarity in our classification system.
4.5.2. Results
Error Outcome Distribution. Table 4 shows the error outcome distribution from the RQ1 experiment on RustRepoTrans, alongside data from Pan et al. (Pan et al., 2024) for translations to other programming languages (C, C++, Python, Java).
In RustRepoTrans, compilation errors account for 94.8% of failures, significantly higher than the 58.3% to 83.3% observed in other benchmarks. This emphasizes the challenge LLMs face when translating to Rust, a language known for stringent compile-time checks due to its ownership model and borrow checker. Unlike other benchmarks, RustRepoTrans recorded no runtime or non-terminating errors, indicating that Rust translations primarily fail at compile-time, underscoring its strong emphasis on memory safety.
The number of compilation errors in unsuccessful translations ranged from 1 to 193, with an average of 9.1 and a median of 4, which complicates debugging. In contrast to languages like Python, where the compiler stops at the first error, the Rust compiler reports all errors, making issue resolution more challenging. We categorized the types of compilation errors based on error codes from (error codes, 2024). Table 5 presents the top 10 compilation errors encountered, with the five most common errors occurring over 1,000 times each. Many of these stem from using non-existent, unimplemented, unresolved, or undeclared elements, highlighting the hallucination phenomenon also seen in other repository-level code generation tasks (Luo et al., 2024; Wang et al., 2024).
Table 4. The comparison of the proportions of error types between RustRepoTrans and (Pan et al., 2024).CE: Compilation Errors, RE: Runtime Errors, FE: Functional Errors, NTE: Non-terminating Execution
Source Target CE RE FE NTE
Pan et al. (Pan et al., 2024) Java, Python, C C++ 74.6% 2.0% 22.0% 1.3%
C++, Python, C Java 74.1% 13.3% 12.4% 0.2%
Java, C++, C Python 58.3% 24.9% 16.4% 0.4%
C++, Java, Python C 83.3% 0.7% 15.5% 0.6%
RustRepoTrans C, Java, Python Rust 94.8% 0% 5.2% 0%
Table 5. Top 10 Compilation Errors Reported by the Rust Compiler on RustRepoTrans
Error Code Frequency Description
E0425 5,949 An unresolved name was used.
E0599 4,601
This error occurs when a method is used on a type
which doesn’t implement it.
E0432 3,504 An import was unresolved.
E0277 2,540
You tried to use a type which doesn’t implement
some trait in a place which expected that trait.
E0433 1,345 An undeclared crate, module, or type was used.
E0609 796 Attempted to access a nonexistent field in a struct.
E0308 744 Expected type did not match the received type.
E0061 222
An invalid number of arguments was passed
when calling a function.
E0412 174 A used type name is not in scope.
E0282 79
The compiler could not infer a type and asked for
a type annotation.
Error Cause Taxonomy. Fig. 6 illustrates the ten error cause categories we identified for failed code translations on RustRepoTrans, specifically those with compilation errors. We further classified these into three main types:
  • Failing to Understand Target Language Features. This type includes errors stemming from the LLM’s insufficient grasp of the syntax and semantics of the target programming language. It encompasses limitations in understanding data types, variable states, and contextual information.
  • Failing to Understand Differences Between Languages. This type relates to the LLM’s inadequate comprehension of the distinctions between the source and target programming languages. It covers issues such as syntax variations, differences in functions, variables, data types, and import paths.
  • Others. This type includes errors unrelated to the LLM’s code translation capabilities. Examples include missing punctuation marks and translations that do not adhere to the provided function signature.
Figure 6. Error Cause Taxonomy for Failed Code Translations on RustRepoTrans in LLMs
The detailed 10 error causes from the types are as following.
Data Type Misinterpretation. This error cause reflects the LLM’s limitations in type inference, particularly evident in Rust’s strict type system. Unlike Python, where variables can change types freely, or C, which allows implicit conversions, Rust requires exact type matches, making translation errors more prominent. LLMs may attempt incompatible assignments, perform arithmetic on mismatched types, or incorrectly treat non-collection variables (like integers) as if they were indexable. In the example below, the model interprets imap_connected_here as a boolean, causing a type mismatch that Rust’s compiler disallows. This illustrates how Rust’s strict typing exposes LLMs’ type inference limitations, often masked in more permissive languages like Python.
let mut imap_connected_here = 0; if imap_connected_here {       # error: expected bool, found integer
Variable State Misinterpretation. This error cause stems from the LLM’s misunderstanding of variable states, leading to several misinterpretations. Common issues include treating undeclared variables as declared, making simultaneous mutable borrows, using the same variable as both mutable and immutable, and incorrectly treating private variables as public. In the example below, the model attempts to pass h as both an immutable reference (&h) and a mutable reference (&mut h), which Rust’s strict borrowing rules prohibit, resulting in a compilation error. These instances highlight the LLM’s limitations in tracking variable states and usage contexts, often leading to significant translation errors.
fn hashit(sha: usize, n: usize, id: &[u8], w: &mut [u8]) -> bool    # hashit function definition hashit(sha, date, &h, &mut h);   # error: using h as both mutable and immutable parameter
Context Misinterpretation. This error cause arises from the LLM’s misunderstanding of context. It includes issues such as failing to recognize when code is within a struct’s built-in function, using asynchronous operations in a synchronous block, or attempting modifications inside a function without the mutable modifier. In the example below, the model incorrectly assumes that batch is a built-in function of a struct. The error occurs when it tries to access self.left(batch.clone()) without using the keyword self, which is necessary in this context.
fn batch(batch: &RecordBatch) -> Result<RecordBatch> { # not a struct’s built-in function let left = self.left(batch.clone())?; # error: use the key word self
Syntactic Differences Misinterpretation. This error cause relates to the LLM’s understanding of syntactic differences between programming languages. It occurs when the model incorrectly assumes that syntax from the source language is valid in the target language. Examples include using chained assignments (allowed in Python, C, and Java) in Rust, attempting to use the goto keyword from C, or employing an unsupported ternary operator ‘‘condition ? expr1 : expr2’’ in Rust.
Function Differences Misinterpretation. This error cause pertains to the LLM’s understanding of functional dependency differences between programming languages. It arises when the model incorrectly assumes that callable functions in the target language are identical to those in the source language. Common issues include assuming a function present in the source language also exists in the target language, that it shares the same scope, or that the invocation method is identical. In the example below, the first line results in an error because the LLM fails to recognize that get should be called within the stock_str namespace.
ret.text = get(context, StockMessage::NoMessages)?;       # error ret.text = stock_str::get(context, StockMessage::NoMessages)?;       # correct
Variable Differences Misinterpretation. This error cause arises from the LLM’s understanding of variable dependency differences between programming languages. It involves errors where the model incorrectly assumes that variables in the target language mirror those in the source language. Common issues include presuming that parameters or declared variables from the source language exist in the target language, that their scopes are the same, or that specific member variables exist within a struct.
Data Type Differences Misinterpretation. This error cause arises from the LLM’s understanding of differences in available data types across programming languages. It reflects errors where the model mistakenly assumes that data types in the source language are also present in the target language. For instance, it may assume that a user-defined data type retains the same name and meaning in both languages.
Dependency Resolution Differences Misinterpretation. This error cause relates to the LLM’s understanding of differences in dependency resolution across programming languages. It highlights errors where the model incorrectly assumes that a specific import path is valid in the target language, such as presuming the existence of a particular dependency library or that certain functions, data types, or elements are present within the imported library.
Missing Punctuation Marks. This error cause arises from the LLM’s failure to include necessary punctuation in its generated output. Examples include incomplete parentheses, missing commas between parameters, and other critical punctuation omissions.
if dc_param_exists(msg->param, DC_PARAM_SET_LATITUDE)) { # error: Redundant closing parenthesis
Function Signature Inconsistencies. This error cause arises from the LLM’s failure to follow instructions to translate according to the provided function signature. Examples include inconsistencies in function modifiers, mismatches in parameter lists, or discrepancies in return values.
Overall Distribution of Error Causes. As shown in Fig. 7, the most common error cause encountered by LLMs during the code translation task on RustRepoTrans is Failing to Understand Differences Between Languages, which accounts for 77.9% of the errors. This is followed by Failing to Understand Target Language Features at 18.5%, and Others at 3.7%. These results indicate that the most significant challenge for LLMs in code translation is effectively grasping the various differences between the source and target languages, such as dependency and syntax differences.
Figure 7. Overall Distribution of Error Causes
At a more detailed level, the top three error causes are Function Differences Misinterpretation (36.6%), Variable Differences Misinterpretation (25.3%), and Data Type Misinterpretation (14.6%). This distribution highlights that function and variable dependencies constitute the largest portion of the dependencies involved in RustRepoTrans code translation tasks. Furthermore, since RustRepoTrans targets a strongly typed language, errors related to data types are ranked third.
Figure 8. Comparative Distribution of Error Causes in LLMs
Comparison of Error Causes Distribution across Different LLMs. Fig. 8 shows the distribution of error causes across various LLMs during code translation on RustRepoTrans. The Failing to Understand Target Language Features error causes are relatively evenly distributed among LLMs due to their similar training data, primarily derived from publicly available code. This results in comparable understanding of the target language across models. In contrast, Failing to Understand Differences Between Languages and Other errors are more common in LLMs with weaker translation capabilities. The former evaluates the model’s understanding of language differences, while the latter reflects its ability to follow instructions, both positively correlating with translation performance.
Notably, Claude-3.5, the top-performing model on RustRepoTrans, exhibits the lowest proportions of Failing to Understand Differences Between Languages and Other error causes. For simpler error types like Syntactic Differences Misinterpretation, Dependency Resolution Differences Misinterpretation, Missing Punctuation Marks, and Function Signature Inconsistencies, the proportions were as low as 2.8%, 2.5%, 0%, and 0%, respectively, indicating superior understanding and adherence to instructions compared to other models.
4.5.3. Conclusion
Existing LLMs struggle with code translation tasks involving dependencies, particularly in distinguishing function and variable differences between source and target languages. When the target language is strongly typed, LLMs often lack understanding of data types. Their ability to recognize language differences is positively correlated with translation performance.
4.6. RQ5: Key Capabilities
Beyond Pass@1 in RQ1, we analyze LLMs’ translation capabilities in noise robustness, syntactical difference identification, and code simplicity.
4.6.1. Design
The key abilities for these three aspects were tested by evaluating each model on a dataset extracted or constructed from RustRepoTrans, tailored to the specific capability being assessed. The same translation prompts used in Section 4.2 were applied. The experimental design for each aspect is described as follows.
Noise Robustness. This capability evaluates the model’s capacity to identify necessary dependencies from provided options, which is crucial in real-world scenarios where dependency information is often uncertain or incomplete. We assessed this capability through two angles: redundancy and incompleteness. For redundancy, we created a dataset by selecting functions and data types with high text similarity (using BLEU scores (Papineni et al., 2002)) relative to the target function’s dependencies, excluding those in the original set. We introduce a novel metric, Redundancy Impact Rate (RIR), to measure the ratio of successful translations between scenarios with Redundant Dependencies and All Dependencies. For incompleteness, we created various datasets by randomly reducing the target function’s dependencies to 75%, 50%, 25%, and 0% of their original size. Using these new benchmarks, we assessed the success rates (Pass@1) of the LLMs and compared them to the original Pass@1 with all dependencies (see RQ1). We introduce a novel metric, Incompleteness Impact Rate (IIR), which measures the LLM’s performance by calculating the average of the Pass@1 under 100%(i.e. All), 75%, 50%, 25%, and 0% dependencies.
Syntactical Differences Identification. This capability assesses the model’s skill in recognizing syntactical differences between source and target programming languages, a key requirement for accurate code translation. For instance, Rust, as a strongly-typed language, does not require checks such as memory exception handling (common in C), type checks (in Python), or null pointer checks (in C and Java) shown in Fig. 9. We selected 89 function pairs from RustRepoTrans that include these checks in the source language and evaluated whether the LLM correctly omits them in the target language. We introduce a novel metric, Syntactical Differences Identification Rate (SDIR@K), which measures the LLM’s success in identifying and omitting these checks in the translated code across $K$ samples.
Figure 9. Examples of checks from other languages that are not needed in Rust
Code Simplicity. This capability assesses the simplicity of translation results, as simpler code is preferred by developers. We evaluate this by calculating token counts with tree-sitter (tree sitter, 2023) and measuring cyclomatic complexity (Gill and Kemerer, 1991), where higher values indicate greater complexity. We apply rust-code-analysis (Ardito et al., 2020) to both the reference and translated code that passed the test cases. A higher ratio suggests that the translated code is more concise and closely matches the reference. We introduce two metrics for code simplicity: Token Rate and CC Rate. These compare the token counts and cyclomatic complexities of the reference and translated code. Higher values for both metrics indicate that the generated code is simpler and more aligned with the original.
Figure 10. The Pass@1 of LLMs under different proportions of dependencies
4.6.2. Results
Noise Robustness. From the perspective of redundant dependencies, after introducing interference dependencies, Claude-3.5 and DeepSeekCoderV2-16B demonstrated minimal fluctuations, with RIR values of 99.4% and 101.6%. These variations fall within the expected range for LLM generation. In contrast, GPT-4 and Llama-3.1-8B exhibited larger fluctuations, with RIR values of 88.2% and 92%, highlighting potential areas for improvement in their noise robustness regarding redundant dependencies. Regarding incomplete dependencies, as shown in Fig. 10, DeepSeekCoderV2-16B and Llama-3.1-8B showed stable performance during gradual reductions, with drop ratios of 19.1% and 18.0% at the 25% level. Noticeable declines occurred only at the Without Dependencies stage. In contrast, Claude-3.5 and GPT-4 faced significant drops at nearly every stage, likely because their successful translations often rely on multiple dependencies. Most successful translations for DeepSeekCoderV2-16B and Llama-3.1-8B came from simpler data with fewer dependencies, ensuring consistent availability across stages except at the Without Dependencies stage. As shown in Fig. 11, the IIR ranking among models are the same as the Pass@1 under All Dependencies in 4.2, Claude-3.5 demonstrated the best performance with IIR values of 32.2%.
Syntactical Differences Identification. As shown in Fig. 11, Claude-3.5 achieves the highest ISDR@1 at 95.5%, demonstrating a strong ability to identify syntactical differences between programming languages. This finding is consistent with the results in Section 4.4 regarding the distribution of error causes across different LLMs (see Fig. 8). In contrast, DeepSeekCoderV2-16B has the lowest ISDR@1 at only 40.5%. This indicates that DeepSeekCoderV2-16B often assumes the checks performed in the source language are also necessary in the target language, leading to errors by incorrectly treating those checks as valid in the translation.
Code Simplicity. Fig. 11 shows that among the successfully translated tasks, the Token Rate for code generated by Claude-3.5 and GPT-4 reached 96.6% and 97.0%, respectively, indicating high simplicity compared to the reference code. In contrast, Llama-3.1-8B had the lowest Token Rate at 63.1%, suggesting it requires more code to achieve the same functionality.
In terms of CC Rate, Claude-3.5 again excels with a CC Rate of 103.7%, while Llama-3.1-8B is notably lower at 53.6%. This difference arises because Claude-3.5 often uses concise built-in functions in place of loops. For instance, as illustrated in Fig. 12, when tasked with finding the first element in an array that meets a condition, Claude-3.5 employs the target language’s any method from the iterator library, whereas other LLMs rely on traditional loops. Furthermore, Claude-3.5’s CC Rate exceeding 1 indicates that, in terms of cyclomatic complexity, its generated code is simpler than the reference code. Fig. 12 shows an example with C as the source language. This simplification likely results from Rust ground truth developers using the C2Rust migration tool (C2Rust, 2024), which translates code through one-to-one pattern matching.
These findings suggest that Claude-3.5 demonstrates a stronger code comprehension and implementation ability, accurately understanding the source code’s functionality and translating it concisely based on the target language’s features.
Figure 11. The performance of LLMs on Key Abilities of code translation.
Figure 12. Claude-3.5 translation result compared to Other LLMs and reference code
4.6.3. Conclusion
The complexity of code generated by LLMs in translation reflects their translation capabilities: models with stronger abilities produce code with lower complexity for equivalent functionality. Claude-3.5, which excels on RustRepoTrans, also shows superior performance in Noise Robustness, Syntactical Differences Identification, and Code Simplicity.
5. Threats to Validity
One potential threat is data leakage between our benchmark and model training data. However, the training data comprises independent function code from different languages rather than functionally equivalent code pairs, which is crucial for code translation. Another concern is the limited size and variety of programming languages in our benchmark, which may impact the generalizability of our findings. We plan to extend our benchmark in the future. Some projects also lack consistent documentation across language versions, leading to potential functional discrepancies. We mitigated this by manually verifying the functionality of the final function pairs. Additionally, the static code analysis tool tree-sitter may not accurately retrieve all dependencies for Rust functions, which we addressed through manual validation. Lastly, our study involved only four models due to time and resource constraints. The focus of this paper is the benchmark itself, not the models. We intend to incorporate more models in future work, but the conclusions and evaluation capabilities of our benchmark remain valid.
6. Conclusion
This work makes the first attempt to evaluate LLMs on repository-level code generation targeting Rust. We first manually construct the first repository-level code translation benchmark RustRepoTrans and perform the first study of 4 state-of-the-art LLMs on repository-level code translation. we find that existing LLMs show much worse performance on repository-level code translation compared to on standalone code translation, LLMs still have significant room for improvement in their code translation capabilities within real-world software development scenarios. Besides, when the target language is a low-resource language with multiple syntactic constraints, such as Rust, LLMs struggle to effectively identify the various differences between the source and target languages, as well as understand the features of the target language.